# Effects of lower floating-point precision on scale-resolving numerical simulations of turbulence

### v1

@@ 8.2 Stagnation
As previously mentioned, the issue of stagnation can become prevalent when using lower precision when summing a large and small number.
The most straightforward and easy to address impact of lower arithmetic precision was observed for example in longer simulations (integration times of $\mathcal{O}(10^{3})$ convective units or larger) with smaller time steps. This was observed with precisions as high as FP32 and as issues with correct estimation of the simulation time (usually calculated as $t_{\rm new}=t_{\rm old}+\Delta t$) as well as some time-dependent if statements (such as those that control the calculation of additional variables at constant intervals). While this was observed mainly in Simson on the longest simulations (integration times of $10000$ convective times or higher) when compiled with FP32, it can easily happen for other codes, especially for smaller time steps. This suggests that integration times $t$ and its related quantities (weights etc.) should use higher precisions (FP64), even when the code is compiled with lower precision.
Another similar observation was the impact of precision on runtime collection of solution statistics. This is related to the addition of the new sample to the previous set, usually done as a variant of $S_{\rm new}=wS_{\rm old}+(1-w)s$, where $S_{\rm new}$ and $S_{\rm old}$ are new and old statistics and $s$ is the new sample. The weighting parameter here, $w=t/(t+\delta t)$, is again highly impacted by precision for short sampling times $\delta t$ (i.e., frequent sampling) and long integration times $t$. This was, for instance, observed as non-zero residuals in the transport equation of Reynolds stresses when calculated from runtime statistics, while not observed when calculated from a few hundred snapshots, despite having far fewer samples.
It was also relevant when computing larger dot-products in Neko, where a naive implementation can run into stagnation issues. This can be remedied through tree-reductions or performing the accumulation in higher precision. This again suggests that to ensure robustness and accuracy of runtime statistics, such calculations and accumulation should be performed in FP64.
@@ 8.3 Arithmetic errors
Among all the computations performed with SSDC 3.1.3, the TS-wave case has shown some implementation details that can be useful in practice. The default FP64 computation is used as a reference, considering the emulated precision with the CPFloat library on both the state and operator terms, reasonable results are obtained as previously discussed. However, when compiling the code in single precision (which is not what the implementation is originally designed for), the computation exhibits nonphysical artifacts in the flow, leading to larger numerical noise and premature transition to turbulence (see Figure 15). Upon close examination, we could trace the origin of the error to the convective term, specifically in the form of a loss of freestream preservation. In simpler terms, the computation of geometric quantities (Jacobians, cell volumes and normals) in single precision resulted in a non-watertight grid. The correct behavior was recovered when switching these computations back to FP64. Similar observations could also be seen when computing the Lambda-2 criterion in Neko, where a naive way of obtaining the required eigenvalues was sensitive to the lower floating point precision. The most straightforward way to remedy this was by making parts of this computation in FP64, as it was neither time nor memory critical.
Figure 15: Evolution in time of the amplitude of 2D modes $|\hat{u}(\alpha,\beta=0)|$ for the transitional case. All simulations are carried out in SSDC. Results from full FP64 and FP32, FP32 emulated with PCS, and compiled FP32 except for the computation of metric terms (MT) in FP64 are shown.
For Simson, it was also evident that in the Chebyshev discretization, points are highly clustered near the ends of the interval, and the derivative matrices are fully populated. As a result, the derivative at any given point is calculated as a weighted sum of values from all other points, with both the weights and function values varying significantly across the domain. A quick examination of the differentiation matrix for first and second derivatives using 257 points (the grid used for the channel flow at $Re_{\tau}\approx 550$) reveals differences of approximately five orders of magnitude between terms in the same row (terms multiplied by values on wall-normal grid points), with the largest discrepancies occurring near the walls. This effect is further amplified in a velocity-vorticity formulation, where higher-order derivatives (up to the fourth order) are required. These factors suggest a high sensitivity to numerical precision, particularly near the wall, which is likely the root cause of the observed issues with the pressure solver in Simson (Section 4.1). Note that the employed Chebyshev-tau method is less sensitive to the poor conditioning.
@@
It is important to highlight that while precision was expected to significantly influence derivatives and flow dynamics, interestingly and somewhat unexpectedly, these sensitivities did not result in irrecoverable errors. In other words, many of the issues were found to be resolved by performing as little as a single time step in FP64.
@@ 9 Conclusions
The main outcome of this work is the demonstrated strong evidence that high-fidelity simulations of wall-bounded turbulence, including direct numerical and large-eddy simulation (DNS and LES), do not necessarily require FP64 arithmetic and can be performed at lower precisions, such as FP32, with minimal impact on the results. We show that this is possible not only for simpler canonical flow cases such as channels, but also for more complex flows exhibiting separation and transition. Our results are consistent across different formulations of governing equations considered in the paper, as well as underlying discretization methods and their implementations in different codes. Essentially perfect overlap is obtained between FP64 and FP32 profiles of various quantities of interest, including, for example, turbulent kinetic energy budgets and high-order statistics. Although we do not focus on performance analysis, we show that using FP32 significant acceleration, up to the ideal 2×2$\times$2 × for a well-optimized code is possible compared to standard FP64. In particular, in FP32 a DNS of selected cases can be run on a single consumer GPU instead of more expensive server-grade cards. In addition, the performance improvement is not limited to computational time but also translates to reduced storage demands. Considering that large-scale CFD simulations utilize hundreds of millions of core hours yearly, all codes should apply significant effort to utilize lower-precision arithmetic, thus saving energy, storage, and money for the same scientific outcomes. However, some caution must be exercised when porting existing solvers, as one likely needs to retain some of the operations in FP64 (see Section 8).
@@
In addition, the paper explores using even lower precisions by means of software emulation. The results are encouraging, with sometimes very low precision, like E5M2, providing good results when applied to only the convective term. A natural way forward is to investigate the native implementation of lower precision formats that are supported by modern hardware. Success is likely conditioned on using mixed precision with the concrete realization of the latter tightly coupled with the numerical method used by the code.
@@
On a subjective note, the results of this study stand in quite strong contradiction with our initial expectations. We anticipated to quickly find a (high) precision threshold, after which the results become unusable. In particular, for flows with transition, we were very skeptical about the possibility of using anything but FP64. However, we were proved wrong and our hope is that this paper will motivate an equally skeptical reader to give reduced precision a chance.
@@
More generally, the question of using arithmetic precisions lower than FP64 should not be viewed as one with a binary yes or no answer, but instead as another hyper-parameter in a simulation. In large-scale simulations, it is often the norm to assess the influence of multiple parameters before the production run starts. These parameters usually include things such as grid resolution, time-step size (i.e. Courant number), domain size, residual thresholds of iterative solvers, and so on. In our view, precision should also become one such parameter, even though the control over it is usually quite rough (i.e., a jump from FP64 to FP32 instead of a smooth transition). In fact, one could argue that even in cases where reduced precision leads to (small) discrepancies in the results, it should still be viewed as just another source of uncertainty in the result, albeit a bias. In that sense, viewing a computational simulation as an optimization problem to maximize the accuracy of the output for a given cost, arithmetic precision must be balanced against other sources of uncertainty such as those related to time averaging, residual, or resolution. Similarly, in parametric studies and simulation campaigns a lower cost per simulation would enable additional simulations, which again should be balanced again other parameters for an objective function of maximizing the knowledge gained from the campaign for a given computational cost.
@@
Despite the wider scope of this study compared to the available literature, it is important to keep in mind that it is still fairly limited compared to the wider applications of high-fidelity simulations. For example, Reynolds numbers of this study were still relatively low and complex geometries or physics were avoided. Future extensions of this work could benefit from a wider variety of test cases, including supersonic and hypersonic flows, combustion, and complex geometries, among other aspects.
@@ Acknowledgments
This project was initiated during the CTR Summer Program 2025 in Stanford [32], for which we gratefully acknowledge the financial support to be able to participate. Computer time was provided by the National Infrastructure for Computing in Sweden (NAISS). The authors gratefully acknowledge the HPC resources provided for this collaborative effort by the Supercomputing Laboratory at King Abdullah University of Science & Technology (KAUST) in Thuwal, Saudi Arabia. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU). The hardware is funded by the German Research Foundation (DFG). This project has received funding from the European High-Performance Computing Joint Undertaking (JU) under grant agreement No 101092621. The JU receives support from the European Union’s Horizon Europe research and innovation programme and Germany, Italy, Slovenia, Spain, Sweden and France.

### v2

@@ 8.2 Stagnation
As previously mentioned, the issue of stagnation can become prevalent when using lower precision when summing large and small numbers.
The most straightforward and easy to address impact of lower arithmetic precision was observed, for example, in longer simulations (integration times of $\mathcal{O}(10^{3})$ convective units or larger) with smaller time steps. This was observed with precisions as high as FP32 and as issues with the correct estimation of the simulation time (usually calculated as $t_{\rm new}=t_{\rm old}+\Delta t$) as well as some time-dependent if statements (such as those that control the calculation of additional variables at constant intervals). While this was observed mainly in Simson on the longest simulations (integration times of $10000$ convective times or higher) when compiled with FP32, it can easily occur for other codes, especially for smaller time steps. This suggests that the integration time $t$ and its related quantities (weights etc.) should use higher precisions (FP64), even when the code is compiled with lower precision.
Another similar observation was the impact of precision on the runtime collection of solution statistics. This is related to the addition of the new sample to the previous set, usually done as a variant of $S_{\rm new}=wS_{\rm old}+(1-w)s$, where $S_{\rm new}$ and $S_{\rm old}$ are new and old statistics and $s$ is the new sample. The weighting parameter here, $w=t/(t+\delta t)$, is again highly impacted by precision for short sampling times $\delta t$ (i.e., frequent sampling) and long integration times $t$. This was, for instance, observed as non-zero residuals in the transport equation of Reynolds stresses when calculated from runtime statistics, while not observed when calculated from a few hundred snapshots, despite having far fewer samples.
It was also relevant when computing larger dot products in Neko, where a naive implementation can run into stagnation issues. This can be remedied through different approaches such as blocked reductions, tree-reductions, or performing the accumulation in higher precision [60, 24]. This again suggests that to ensure robustness and accuracy of runtime statistics, such calculations and accumulation should be performed in FP64.
@@ 8.3 Arithmetic errors
Among all the computations performed with SSDC 3.1.3, the TS-wave case has shown some implementation details that can be useful in practice. The default FP64 computation is used as a reference, considering the emulated precision with the CPFloat library on both the state and operator terms, reasonable results are obtained as previously discussed. However, when compiling the code in single precision (which is not what the implementation is originally designed for), the computation exhibits nonphysical artifacts in the flow, leading to larger numerical noise and premature transition to turbulence (see Figure 15). Upon close examination, we could trace the origin of the error to the convective term, specifically in the form of a loss of freestream preservation. In simpler terms, the computation of geometric quantities (Jacobians, cell volumes, and normals) in single precision resulted in a non-watertight grid. The correct behavior was recovered when these computations were switched back to FP64. Similar observations could also be seen when computing the Lambda-2 criterion in Neko, where a naive method for obtaining the required eigenvalues was sensitive to the lower floating point precision. The most straightforward way to remedy this was by making parts of this computation in FP64, as it was neither time nor memory critical.
Figure 15: Evolution in time of the amplitude of 2D modes $|\hat{u}(\alpha,\beta=0)|$ for the transitional case. All simulations are carried out in SSDC. Results from full FP64 and FP32, FP32 emulated with PCS, and compiled FP32, except for the computation of metric terms (MT) in FP64, are shown.
For Simson, it was also evident that in the Chebyshev discretization, points are highly clustered near the ends of the interval, and the derivative matrices are fully populated. As a result, the derivative at any given point is calculated as a weighted sum of values from all other points, with both the weights and function values varying significantly across the domain. A quick examination of the differentiation matrix for first and second derivatives using 257 points (the grid used for the channel flow at $Re_{\tau}\approx 550$) reveals differences of approximately five orders of magnitude between terms in the same row (terms multiplied by values on wall-normal grid points), with the largest discrepancies occurring near the walls. This effect is further amplified in a velocity-vorticity formulation, where higher-order derivatives (up to the fourth order) are required. These factors suggest a high sensitivity to numerical precision, particularly near the wall, which is likely the root cause of the observed issues with the pressure solver in Simson (Section 4.1). Note that the employed Chebyshev-tau method is less sensitive to poor conditioning.
@@
It is important to highlight that while precision was expected to influence derivatives and flow dynamics significantly, interestingly and somewhat unexpectedly, these sensitivities did not result in irrecoverable errors. In other words, many of the issues were found to be resolved by performing as little as a single time step in FP64.
@@ 9 Conclusions
The main outcome of this work is the demonstrated strong evidence that high-fidelity simulations of wall-bounded turbulence, including direct numerical and large-eddy simulation (DNS and LES), do not necessarily require FP64 arithmetic and can be performed at lower precisions, such as FP32, with minimal impact on the results. We demonstrated that this is possible not only for simpler canonical flow cases such as channels, but also for more complex flows that exhibit separation and transition. Our results are consistent across different formulations of governing equations considered in the paper, as well as underlying discretization methods and their implementations in different codes. Perfect overlap is obtained between FP64 and FP32 profiles of various quantities of interest, including, for example, turbulent kinetic energy budgets and high-order statistics. Although we do not focus on performance analysis, we show that using FP32, significant acceleration, up to the ideal $2\times$ for a well-optimized code, is possible compared to standard FP64. In particular, in FP32, a DNS of selected cases can be run on a single consumer GPU instead of more expensive server-grade cards. In addition, the performance improvement is not limited to computational time but also translates to reduced storage demands. Considering that large-scale CFD simulations utilize hundreds of millions of core hours yearly, all codes should apply significant effort to utilize lower-precision arithmetic, thus saving energy, storage, and money for the same scientific outcomes. However, some caution must be exercised when porting existing solvers, as one likely needs to retain some of the operations in FP64 or adapt the implementation to make it more robust to arithmetic precision (see Section 8).
@@
In addition, the paper explores the use of even lower precisions through software emulation. The results are encouraging, with sometimes very low precision, like E5M2, providing good results when applied to only the convective term. A natural way forward is to investigate the native implementation of lower precision formats supported by modern hardware. Success is likely conditioned on using mixed precision with the concrete realization of the latter tightly coupled with the numerical method used by the code.
@@
On a subjective note, the results of this study stand in quite strong contradiction to our initial expectations. We anticipated quickly finding a (high) precision threshold, after which the results become unusable. In particular, for flows with transition, we were very skeptical about the possibility of using anything but FP64. However, we were proved wrong, and our hope is that this paper will motivate an equally skeptical reader to give reduced precision a chance.
@@
More generally, the question of using arithmetic precisions lower than FP64 should not be viewed as one with a binary yes or no answer, but instead as another hyper-parameter in a simulation. In large-scale simulations, it is often the norm to assess the influence of multiple parameters before the production simulation starts. These parameters usually include things such as grid resolution, time-step size (i.e. Courant number), domain size, residual thresholds of iterative solvers, and so on. In our view, precision should also become one such parameter, even though the control over it is usually quite rough (i.e., a jump from FP64 to FP32 instead of a smooth transition). In fact, one could argue that even in cases where reduced precision leads to (small) discrepancies in the results, it should still be viewed as just another source of uncertainty in the result, albeit a bias. In that sense, viewing a computational simulation as an optimization problem to maximize the accuracy of the output for a given cost, arithmetic precision must be balanced against other sources of uncertainty such as those related to time averaging, residual, or resolution. Similarly, in parametric studies and simulation campaigns a lower cost per simulation would enable additional simulations, which again should be balanced against other parameters for an objective function of maximizing the knowledge gained from the campaign for a given computational cost.
@@
Despite the broader scope of this study compared to the available literature, it is important to keep in mind that it is still fairly limited compared to the wider applications of high-fidelity simulations. For example, the Reynolds numbers of this study were still relatively low, and complex geometries or physics were avoided. Future extensions of this work could benefit from a wider variety of test cases, including supersonic and hypersonic flows, combustion, and complex geometries, among other aspects.
@@ Acknowledgments
This project was initiated during the CTR Summer Program 2024 in Stanford [34], for which we gratefully acknowledge the financial support that enabled our participation. Computer time was provided by the National Infrastructure for Computing in Sweden (NAISS). The authors gratefully acknowledge the HPC resources provided for this collaborative effort by the Supercomputing Laboratory at King Abdullah University of Science & Technology (KAUST) in Thuwal, Saudi Arabia. This project has received funding from KAUST under grant No. BAS/1/1663-01-01. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU). The hardware is funded by the German Research Foundation (DFG). This project has received funding from the European High-Performance Computing Joint Undertaking (JU) under grant agreement No. 101092621. The JU receives support from the European Union’s Horizon Europe research and innovation programme and Germany, Italy, Slovenia, Spain, Sweden and France.